Suppose TC = 480 + 10Q + 4Q^2. When Q = 10. What is Fixed Cost and Variable Cost?
1
Expert's answer
2020-03-26T09:31:56-0400

If "TC = 480 + 10Q + 4Q^2" , and Q = 10, then:

Fixed Cost FC = 480,

Variable Cost "VC = 10Q + 4Q^2 = 10\u00d710 + 4\u00d710^2 = 500."